OpenStudy (literallyjian):

Help! A superhero flies 160 m from the top of a tall building at an angle of 20 ◦ below the horizontal. What is the horizontal component of the superhero’s displacement? Draw the vectors to scale on a graph to determine the answer. Answer in units of m. Your answer must be within ± 5.0% What is the vertical component of the superhero’s displacement? Answer in units of m. Your answer must be within ± 5.0% I've gotten the first part, it's 150.612 but I can't manage to get the first part, whether it's Pythagoras or sin/cos. 54.anything is not a right answer.

OpenStudy (sooobored):

@Directrix the solution looks correct, am i missing anything?

Directrix (directrix):

Look at the attachment and let me know if it makes sense. @sooobored

OpenStudy (literallyjian):

Y is 150.612 the problem is that it says 54.723 is wrong

Directrix (directrix):

Vertical Component https://sehsphysics.wikispaces.com/file/view/summer+quest+2+KEY.pdf

OpenStudy (literallyjian):

I've gotten 54.723 so many times I want to screaM but it says that it's not correct!

OpenStudy (austin.l):

We start out with this correct? 

OpenStudy (austin.l):



OpenStudy (literallyjian):

no we only start with the hypotenuse and the angle °

OpenStudy (literallyjian):

yeah that looks right :)

OpenStudy (austin.l):

We can then find x by using cos. \[\cos{(20)} = \frac{x}{160}\] Solving for x, we find x = 150.35 We can then find y using the pythagorean theorem. \[a^2 + b^2 = c^2\]

OpenStudy (literallyjian):

just the things ive tried that were all wrong

Directrix (directrix):

@Austin.L Should the equation be this: cos (25) = x/160 ?

OpenStudy (literallyjian):

the angle° is 20

OpenStudy (austin.l):

Or rather, \[x^2 + y^2 = h^2\] We know, h = 160 & x = 150.35 \[(150.35)^2 + y^2 = (160)^2\] Solving for y we find y = 54.7255

Directrix (directrix):

Okay, the attachments will give an example of how to work the problem but the angle of the attachment problem is not the same as that of this problem. Sorry to mislead.

OpenStudy (literallyjian):

it's no problem! :)

OpenStudy (austin.l):

Or we can go about it using trig. \[\sin{(20)} = \frac{x}{160}\] Which when you solve for x is 54.723 I'm not sure why you'd have anything else.

OpenStudy (literallyjian):

that's what I'm saying! but it always counts it as wrong, and im on my last try here so I don't knwo

OpenStudy (austin.l):

I've solved it with trig formulas, and cos and sin... and I get approximately the same answer. I'm not sure why it says its wrong man.

OpenStudy (literallyjian):

ill try it again, maybe glitching! thanks for your help!! :)

OpenStudy (austin.l):

No problem buddy!

OpenStudy (mhchen):

Yeah, I feel there might be something missing... I'd rather have you post a screeny of the question next time.

OpenStudy (literallyjian):

wrong :(

OpenStudy (mhchen):

what is like the format

OpenStudy (mhchen):

of ur answer input, like does the website check it to some sigfig?

OpenStudy (literallyjian):

no sigfig checks

OpenStudy (mhchen):

That looks about right....so you got 150.351m for horizontal and -54.723m for vertical...

OpenStudy (mhchen):

Have you tried inserting a negative in? so instead of 54.7 Make it -54.7

OpenStudy (austin.l):

Wait... waiiiiiiiit.... It would be a negative!!!

OpenStudy (literallyjian):

DARN IT!! I put in my last answer so it won't accept any more!! :S

OpenStudy (mhchen):

Yeah, happens in 2Dimensional problem.s

OpenStudy (mhchen):

oh well, rip 10 points.

OpenStudy (austin.l):

Here we are.

OpenStudy (austin.l):

I'm so sorry, I forgot that it was 20 degrees down which means the displacement would be down meaning negative! D':

OpenStudy (literallyjian):

why would it be negative over normal?

OpenStudy (mhchen):

Actually it's negative cuz it's below the horizontal, that's it.

OpenStudy (mhchen):

If it was 20 degrees above the horizontal, then the vertical component would be positive.

OpenStudy (literallyjian):

darN!